On August 1, 2017, Alpha Company retired a $10,000 bond at 102. On July 31, 2017 Alpha had accrued the interest payment for $250. At the date of retirement, the unamortized discount on the bonds was $400. Use this information to determine the dollar value of Alpha Companys: (Round all dollar values to the nearest whole dollar.) 1. Gain or Loss on Retirement of Bond: enter any loss dollar & amount inside of brackets ) 2. Cash paid to retire the bond issue 3. Amount of Interest Payable liquidated at retirement 4. Amount of Premium or Discount on Bond Payable liquidated at retirement
Answer #1

1.Loss on retirement of bond=

=Amount paid over face value value +Unamortized discount - Accrued interest

​​​​​​ =(10,000*0.02)+400-250

=$350

2.Cash paid to retire the bond issue=10000*102%=$10200

3.Amount of Interest Payable liquidated at retirement=$250

4.Amount of Premium or Discount on Bond Payable liquidated at retirement=$400

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW. PLEASE RATE MY ANSWER BY HITTING ??. THANK YOU